OpenStudy (anonymous):

You invest $2,000.00 at a rate of 4% APR for a 3 months. How much interest did you earn? $80.00 $40.00 $20.00 $10.00 You earn $10.00/hour and work 40 hours/week. You invest $4,000.00 at an interest rate of 5% APR for a year. How much interest did that savings earn in a year in terms of hours of work? 100 hours 20 hours 40 hours 400 hours

OpenStudy (anonymous):

3 months = 1/4 year rate is 4 principle is 2000 interest = PRT/100

OpenStudy (anonymous):

^^ Huh?

OpenStudy (anonymous):

what huh ? lol

OpenStudy (anonymous):

Can u help, explain further. I like explanations.

OpenStudy (anonymous):

time is 3 month.. so it will be 1/4 years right ?

OpenStudy (anonymous):

Well, yes. I got that.

OpenStudy (anonymous):

number one: (4/100)(3/12)(2000.00) number two: (4000.00)(5/100) divide this by 10 for answer in hours

OpenStudy (anonymous):

simple interest = principal*time*rate/100 = 2000*4*(1/4)/100 = $20

OpenStudy (anonymous):

Ohhh okay, I understand that now. So what about number two?

OpenStudy (anonymous):

ok